The cryoscopic constant of water is 1.86.a 0. 0 1 molal Acetic Acid solution produces a depression of. 01 93degree celsius in freezing point. degree of dissociation of Acetic acid is

Grade:12

1 Answers

Patel priyanka
6 years ago
Degree of dissociation=i-1÷n-1Here,n=2∆Tf=0.0193℃, Kf=1.86 K kg/mol,m=0.01m∆Tf=iKfmi=∆Tf÷Kfm =0.0193÷1.86×0.01=1.037Degree of dissociation=1.037-1÷2-1 =0.037